OpenStudy (anonymous):

Find the couterexemple to show that the equation 1-sinx = cos x is not an identity

OpenStudy (anonymous):

identity holds for every value of the variable. put x=pi/4; \[1-\frac{1}{\sqrt2} \neq \frac{1}{\sqrt2}\]

OpenStudy (anonymous):

1 - sinx = cosx 1= sinx +cosx x = pi/4

OpenStudy (anonymous):

ty